The measures of the angles of the isosceles triangle are in the ratio 8:8:2
Answer by stanbon(75887)

The measures of the angles of the isosceles triangle are in the ratio 8:8:2
-----
8:8:2 is the same as 8x:8x:2x
-----
Equation:
8x + 8x + 2x = 180 degrees
18x = 180 degrees
x = 10
------
8x = 8*10 = 80 degrees
2x = 2*10 = 20 degrees
8x = 8*10 = 80 degrees
==========================
Cheers,
Stan H.